Reports 1-1 of 1 Clear search Modify search
fabian.arellano - 17:52 Tuesday 03 September 2019 (10291) Print this report
New optic alignment and vertical OSEMs positions

With Yokozawa-san.

We moved the tilt QPD according to the new optic orientation required by the interferometer.

  1. We did it by keeping the alignmnent of the optic with the oplev DC control off and providing actuation with the OPTIC ALIGN blosk in the IM section.
  2. We checked the interference pattern was still produced after moving the tilt QPD.
  3. Somehow we forgot to define the new set point, so we chose one close to the centre of the QPD which should be close to the required value for the interferometer.
  4. The diagonalization of the oplev should be checked again.
  5. For the new orientation for the ALIGNED state there was a change in IM-P of +626 µrad with respect to the old one. Namely, it went from -140 µrad to 486 µrad. The readout of the vertical OSEMs are:
    • OSEM #59, V1: 5010 cnt, -211 µm.
    • OSEM #60, V2: 3247 cnt, -299 µm.
    • OSEM #61, V3: 3837 cnt, -264 µm.

I took a transfer function in order to assess the health of the vertical OSEMs alignment. It looks alright, however, V2 looks close to the end of the linear range. According to calibration data it's about 140 µm from the end of such a range.  For reference see the calibration curves at klog entries 5914, 6521 and 6460.






Images attached to this report
Comments to this report:
ayaka.shoda - 9:38 Wednesday 04 September 2019 (10315) Print this report

Hi, thanks for the work.

I think it would be good to check the sensor coupling as well.

Please try to use the auto-measurement script. It has the referenses of other DOFs. (and the references should be newer than the one you are using.)
It would not take much time, since I reduced the frequency resolutions and averages.

fabian.arellano - 13:51 Wednesday 04 September 2019 (10318) Print this report

I used the template used in the auto-measurement and modified it as follows:

  • I updated the reference, the old one was from SR2 not SRM.
  • I incresed the resolution from 30 mHz to 10 mHz in order to see the resonant peaks clearly. With 30 mHz the peaks were not as clear.
  • I just noticed the number of A channels is short by a number of three. I will update it and do the measurement again.
  • I will measure transfer functions for TM-P and TM-R also.
  • The cross-coupling between the different degrees of freedom of all the IMs in the Type B suspensions still have to be systematically assessed. A procedure should be written.

It's worth pointing out that before setting the new orientation of the optic the V2 OSEM was already in a position close to its current one. As reported in entry 9028 from the 31st of May the position fo the OSEM flag was 151 µm (3380 counts) from the end of its linear range. This is related to a problem with the Bottom Filter. We cannot move it lower than -30 µm and it was moved to +100 µm for precaution.

In the current position and orientation of the optic, the IM is 140 µm below the upper limit (defined by the OSEM linear range) and 130 µm above the lower limit (defined by the condition of the BF keystone).

fabian.arellano - 15:17 Wednesday 04 September 2019 (10323) Print this report

In case you don't see the pictures attached to this report please click here: 10323

The masurement of the transfer functions in IM-V, IM-R and IM-P suggest the vertical OSEMs are functional at the current position.

The following applies:

  • The reference used for IM-V is a recent one.
  • The references used for IM-P and IM-R are old ones from SR2. I will update them with the current measurement.
  • Crossed-coulping data for all the degrees of freedom of the IM was measured and is within the file. Analysis is pending.




  • SRM_IMP_exc.xml
  • SRM_IMR_exc.xml
  • SRM_IMV_exc.xml
Images attached to this comment
ayaka.shoda - 9:08 Thursday 05 September 2019 (10342) Print this report

Have you checked the new auto-measurement set? That is in /kagra/Dropbox/Subsystems/VIS/AutoMeasurement/.

The templates there are measurement of SRM for sure, which were measured on July 26.

Search Help